# LLMs, You Can Evaluate It! Design of Multi-perspective Evaluation for Reports in Security Operation Centers

### v2

@@ 7.0.1 Limitations
Our study has several limitations that need to be addressed in future work. First, we utilized private datasets as analysis reports. While pseudo reports are also used to support reproducibility, the number of reports and their length are limited, and hence, we need to extend them into a more diverse setting. Second, the Analyst-wise Checklist contains a large number of items, and the important evaluation items may differ depending on the reports. Future work will concentrate on a mechanism that automatically selects these items, although they were manually selected in this paper. Third, the number of SOC practitioners involved in the human gold evaluation was limited. To confirm the impact of MESSALA on reducing analysis workload and improving the quality of analysis reports in real-world environments, we also need to conduct further investigation with a larger number of participants. Finally, the evaluation perspectives in this paper, including the guidelines and defect categories, are based on the literature review and the semi-structured interviews with practitioners and do not fully cover the evaluation processes and defect patterns across all real-world SOCs. We are in the process of refining feedback from versatile operation contexts.
@@ 8 Conclusion
In this paper, we discussed evaluations of analysis reports in SOCs using LLMs. We first designed the Analyst-wise Checklist by identifying evaluation criteria for analysis reports in SOCs from multiple perspectives of veteran SOC practitioners as the answer to RQ1 through literature review and interviews with SOC practitioners. Next, we proposed MESSALA that can maximize evaluations of analysis reports with feedback comments from the standpoint of human SOC practitioners. We also conducted quantitative evaluations of analysis reports using MESSALA and demonstrated that MESSALA can rate the evaluation criteria for analysis reports, thereby enabling LLMs to quantitatively evaluate the reports compared with the existing methods as the answer to RQ2. Finally, we conducted qualitative evaluations of analysis reports using MESSALA in terms of the feedback comments on the reports. We then showed that MESSALA can qualitatively evaluate analysis reports and provide actionable feedback comments under the evaluation criteria as the answer to RQ3. We plan to conduct further experiments across diverse settings, including larger sample sizes and feedback comments from diverse operation contexts, as well as additional deployments in real-world SOC environments.

### v3

@@ 7.0.1 Limitations
Our study has several limitations that need to be addressed in future work. First, we utilized private datasets as analysis reports. While pseudo-reports are also used to support reproducibility, the number of reports and their length are limited, and hence, we need to extend them into a more diverse setting. Second, the Analyst-wise Checklist contains a large number of items, and the important evaluation items may differ depending on the reports. Future work will concentrate on a mechanism that automatically selects these items, although they were manually selected in this paper. Third, the number of SOC practitioners involved in the human gold evaluation was limited. To confirm the impact of MESSALA on reducing analysis workload and improving the quality of analysis reports in real-world environments, we also need to conduct further investigation with a larger number of participants. Finally, the evaluation perspectives in this paper, including the guidelines and defect categories, are based on the literature review and the semi-structured interviews with practitioners and do not fully cover the evaluation processes and defect patterns across all real-world SOCs. We are in the process of refining feedback from versatile operation contexts.
@@ 8 Conclusion
In this paper, we discussed evaluations of analysis reports in SOCs using LLMs. We first designed the Analyst-wise Checklist by identifying evaluation criteria for analysis reports in SOCs from multiple perspectives of veteran SOC practitioners as the answer to RQ1 through the literature review and interviews with SOC practitioners. Next, we proposed MESSALA, which can maximize the evaluation of analysis reports with feedback comments from the perspective of human SOC practitioners. We also conducted quantitative evaluations of analysis reports using MESSALA and demonstrated that MESSALA can rate the evaluation criteria for analysis reports, thereby enabling LLMs to quantitatively evaluate the reports compared with the existing methods as the answer to RQ2. Finally, we conducted qualitative evaluations of analysis reports using MESSALA in terms of the feedback comments on the reports. We then showed that MESSALA can qualitatively evaluate analysis reports and provide actionable feedback comments under the evaluation criteria as the answer to RQ3. We plan to conduct further experiments across diverse settings, including larger sample sizes and feedback comments from diverse operation contexts, as well as additional deployments in real-world SOC environments.